Ticket Number: HEA-NOC/19981218-01 Ticket Status: INFORMATIONAL
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: HEAnet <==> ITnet
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: 19981218 17:52:14 UTC Problem Start: 19981217 14:41 UTC
Ticket Closed: 19981218 17:52:14 UTC Problem End: 19981217 14:42 UTC
Emergency change of ITnet's connection at the NOC to another
ITnet was disconnected for approx 20 seconds (2 seconds for the
physical change, the rest while routing reestablished itself).
Checked change with Lancomms before proceeding; configured new
interface, switched to G.703 interface. NOC are continuing to monitor;
all looks well.
Time to Fix: